What is the answer to: 8/4+5r/4=6r/2 I have tried everything I know but I still can't get it.
Answer by Alan3354(69443)

What is the answer to: 8/4+5r/4=6r/2 I have tried everything I know but I still can't get it.
---------------
What is the question?
If you want to solve for r:
8/4 + 5r/4 = 6r/2
2 + 5r/4 = 3r
8 + 5r = 24r
19r = 8
r = 8/19